Solve the following system of linear equations
x+ y- 3z+ 2w = 0
2x -y +2z- 3w =0
3x- 2y+ z -4w =0
-4x +y -3z +w =0
1
Expert's answer
2020-06-04T19:35:48-0400
"A=\begin{pmatrix}
 1 & 1 & -3 & 2 \\
 2 & -1 & 2 & -3 \\
 3 & -2 & 1 & -4 \\
 -4 & 1 & -3 & 1
\end{pmatrix}, X=\begin{pmatrix}
 x \\
 y \\
 z \\
 w 
\end{pmatrix}, B=\begin{pmatrix}
 0 \\
 0 \\
 0 \\
 0 
\end{pmatrix}"

"AX=B"

"R_2=R_2-(2)R_1"


"\begin{pmatrix}
 1 & 1 & -3 & 2 \\
 0 & -3 & 8 & -7 \\
 3 & -2 & 1 & -4 \\
 -4 & 1 & -3 & 1
\end{pmatrix}"

"R_3=R_3-(3)R_1"


"\begin{pmatrix}
 1 & 1 & -3 & 2 \\
 0 & -3 & 8 & -7 \\
 0 & -5 & 10 & -10 \\
 -4 & 1 & -3 & 1
\end{pmatrix}"

"R_4=R_4+(4)R_1"


"\begin{pmatrix}
 1 & 1 & -3 & 2 \\
 0 & -3 & 8 & -7 \\
 0 & -5 & 10 & -10 \\
 0 & 5 & -15 & 9
\end{pmatrix}"

"R_2=R_2\/(-3)"


"\begin{pmatrix}
 1 & 1 & -3 & 2 \\
 0 &1 & -8/3 & 7/3 \\
 0 & -5 & 10 & -10 \\
 0 & 5 & -15 & 9
\end{pmatrix}"

"R_1=R_1-R_2"


"\begin{pmatrix}
 1 & 0 & -1/3 & -1/3 \\
 0 & 1 & -8/3 & 7/3 \\
 0 & -5 & 10 & -10 \\
 0 & 5 & -15 & 9
\end{pmatrix}"

"R_3=R_3+(5)R_2"


"\begin{pmatrix}
 1 & 0 & -1/3 & -1/3 \\
 0 & 1 & -8/3 & 7/3 \\
 0 & 0 & -10/3 & 5/3 \\
 0 & 5 & -15 & 9
\end{pmatrix}"

"R_4=R_4-(5)R_2"


"\begin{pmatrix}
 1 & 0 & -1/3 & -1/3 \\
 0 & 1 & -8/3 & 7/3 \\
 0 & 0 & -10/3 & 5/3 \\
 0 & 0 & -5/3 & -8/3
\end{pmatrix}"

"R_3=(-3\/10)R_3"


"\begin{pmatrix}
 1 & 0 & -1/3 & -1/3 \\
 0 & 1 & -8/3 & 7/3 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & -5/3 & -8/3
\end{pmatrix}"

"R_1=R_1+(1\/3)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & -1/2 \\
 0 & 1 & -8/3 & 7/3 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & -5/3 & -8/3
\end{pmatrix}"

"R_2=R_2+(8\/3)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & -1/2 \\
 0 & 1 & 0 & 1 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & -5/3 & -8/3
\end{pmatrix}"

"R_4=R_4+(5\/3)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & -1/2 \\
 0 & 1 & 0 & 1 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & 0 & -7/2
\end{pmatrix}"

"R_4=(-2\/7)R_4"


"\begin{pmatrix}
 1 & 0 & 0 & -1/2 \\
 0 & 1 & 0 & 1 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & 0 & 1
\end{pmatrix}"

"R_1=R_1+(1\/2)R_4"


"\begin{pmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & 0 & 1
\end{pmatrix}"

"R_2=R_2-R_4"


"\begin{pmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & -1/2 \\
 0 & 0 & 0 & 1
\end{pmatrix}"

"R_3+R_3+(1\/2)R_4"


"\begin{pmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1
\end{pmatrix}"

"\begin{pmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
 z \\
 w 
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0 \\
 0 
\end{pmatrix}"

"x=0"

"y=0"

"z=0"

"w=0"


"(0, 0,0,0)"